OpenStudy (anonymous):

Find the probability of choosing a point from the shaded region given below. The radius of the inner circle is 3 inches and the radius of the outer circle is 9 inches.

OpenStudy (anonymous):

what kind of question is this....there are infinitely many possibilities.

OpenStudy (anonymous):

you cannot have a finite set of points within a set of area.

OpenStudy (anonymous):

A. 14/127 B.1/4 C.3/6 D.1/9

OpenStudy (anonymous):

OH oops im sorry, I misunderstood

OpenStudy (anonymous):

do you know how to find area of a circle? the probability is the area of the little circle divided by the area of the little circle.

OpenStudy (anonymous):

no , can you help me find it? please @sylbot

OpenStudy (anonymous):

area of a circle with respect to radius is \[\pi \times r^2\]

OpenStudy (anonymous):

so just plug in the radii for each circle and take the area of the little one over the area of the big one to find the probability.

OpenStudy (anonymous):

is the radius 8?

OpenStudy (anonymous):

@sylbot?

OpenStudy (anonymous):

nope, you said it yourself in the question. The radius is 9 and 3 for the bigger and smaller circles respectively.

OpenStudy (anonymous):

pi * 9^2 = 254.46 =254.5 pi * 3^2 = 28.27 = 28.3 so what now? @sylbot

OpenStudy (anonymous):

now the probability is the smaller area divided by the bigger area.

OpenStudy (anonymous):

0.111198428 @sylbot?

OpenStudy (anonymous):

sounds right to me!

OpenStudy (accessdenied):

You might want to leave it like this: \( \dfrac{ \pi \times 3^2 }{ \pi \times 9^2 } \) And simplify the fraction by cancelling like factors in numerator and denominator: \( \dfrac{ \pi \times 9 }{\pi \times 9 \times 9 } \) That will give an answer from the answer choices. Otherwise just go through each choice and figure the fraction in decimal form to compare.

OpenStudy (anonymous):

1/9 . its D ! thanks again @AccessDenied

OpenStudy (anonymous):

thanks @sylbot